OpenStudy (anonymous):

write the equation of the line through each point. use slope intercept form. through (7,1) and perpendicular to y=-x+3.

OpenStudy (anonymous):

y = -x + 3 do you know what the slope is in this equation ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

y = mx + b....m is your slope, therefore the slope to that equation is -1. Understand so far ?

OpenStudy (anonymous):

ok got it.

OpenStudy (anonymous):

We are looking for a perpendicular line, therefore the slope will change. We want the negative reciprocal of slope -1. That just means " flip " the slope and change the sign. Therefore, the slope we will need will be 1. Understand ?

OpenStudy (anonymous):

yes ok.

OpenStudy (anonymous):

We will now use point slope formula : y - y1 = m(x - x1) using slope(m) = 1 and (7,1) x1 = 7 and y1 = 1 so lets sub our info into the formula.. y - 7 = 1(x - 1) you still with me ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

y - 7 = 1(x - 1) y - 7 = x - 1 y = x - 1 + 7 y = x + 6 any questions ?

OpenStudy (anonymous):

thx and i have no questions.

OpenStudy (anonymous):

if you were looking for a parallel line, the slope would have stayed the same. But perpendicular is negative reciprocal of the slope. Remember this.

OpenStudy (anonymous):

ok thx again.

OpenStudy (anonymous):

glad to help :)